OpenStudy (anonymous):

whats 500 divided by 60

OpenStudy (anonymous):

can you factor?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[ 500=2 \times250=2\times2\times125=2\times2\times5\times25=2\times2\times5\times5\times5 \]

OpenStudy (anonymous):

\[ 60=2\times 30=2\times 2\times 15=2\times2\times3\times5 \]

OpenStudy (anonymous):

Cancel out common factors

OpenStudy (anonymous):

so 20?

OpenStudy (anonymous):

\[ \frac{2\times2\times5\times5\times5}{2\times2\times3\times5} = \frac{\cancel2\times\cancel2\times\cancel5\times5\times5}{\cancel2\times\cancel2\times3\times\cancel5} = \frac{5\times5}{3} = \frac {25}3 \]

OpenStudy (anonymous):

That is the most reduced form

OpenStudy (anonymous):

thanks!!1